Please help me with it
What was the weight percent of water in the hydrat
Answer #1

2. Weight percentage of water before heating = (weight of water/weight of hydrate)*100 = (0.505/2.029)*100 = 24.88%

3. NUmber of moles of water = weight of H2O/mol. wt of H2O = 0.415/18 = 0.023 moles

4. The given formula contain 1mole of H2O for eacch mole of formula of compound

5. Weight of crucible = 3.715 g

Weight of hydrate added = 2.000 g

Total weight = 3.715+2.000 = 5.715 g

Weight of crucible+anhydrous salt = 5.168

Weight of water = 5.715 + 5.168 = 0.547 g